OpenStudy (anonymous):

Where does the graph of y = 6x2 + 5x − 6 cross the x-axis? (−1/6, 0) and (6, 0) (1/6, 0) and (−6, 0) (2/3, 0) and (−3/2, 0) (−2/3, 0) and (3/2, 0)

OpenStudy (unklerhaukus):

0 = 6x2 + 5x − 6

Parth (parthkohli):

Hint: The \(\mathsf{x-intercept}\).

Parth (parthkohli):

And then I am too lazy to solve it for you. Give it a try.

OpenStudy (anonymous):

they 6x2 is 6x^2

OpenStudy (anonymous):

the*

Parth (parthkohli):

Yep.

OpenStudy (anonymous):

parth i cant believe u

OpenStudy (precal):

Partkohli is correct, you are looking for the x intercepts, also known as roots or solutions

OpenStudy (precal):

if it is factorable, you can factor it to solve it. you can always use the quadratic formula to find the solutions as well (it works 100% of the time) you can also, look at the graph to determine the solutions. Many ways to do this problem

OpenStudy (unklerhaukus):

\[0 = ax^2 + bx +c\]\[\Rightarrow\qquad x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[0 = 6x^2 + 5x − 6\]\[\Rightarrow\qquad x=\frac{-(5)\pm\sqrt{(5)^2-4\times(6)\times(-6)}}{2\times(6)}\]

OpenStudy (anonymous):

ok thanks i can solve that

OpenStudy (unklerhaukus):

to get from \[y= 6x^2 + 5x − 6\]to \[0 = 6x^2 + 5x − 6\] i have used the fact that the \(x\)-intercept is when y=0